OpenStudy (anonymous):

Use any model you choose to solve the story problem. A single board is 1/12 meter thick. How many boards are there in a stack of boards that is 2 meters high?

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

you just need to divide 2 by 1/12 \[\frac{ 2 }{ 1/12 }=\frac{2*12}{1/12*12}=24\]

OpenStudy (anonymous):

24 boards